### v2

# Pretraining and Updating Language- and Domain-specific Large Language Model: A Case Study in Japanese Business Domain

Abstract
Several previous studies have considered language- and domain-specific large language models (LLMs) as separate topics. This study explores the combination of a non-English language and a high-demand industry domain, focusing on a Japanese business-specific LLM. This type of a model requires expertise in the business domain, strong language skills, and regular updates of its knowledge. We trained a 13-billion-parameter LLM from scratch using a new dataset of business texts and patents, and continually pretrained it with the latest business documents. Further we propose a new benchmark for Japanese business domain question answering (QA) and evaluate our models on it. The results show that our pretrained model improves QA accuracy without losing general knowledge, and that continual pretraining enhances adaptation to new information. Our pretrained model and business domain benchmark are publicly available.  [^1] https://huggingface.co/stockmark
1 Introduction
This study explores a language- and domain-specific large language model (LLM), a Japanese LLM for the business domain. As examples of business-related questions shown in Table 1, answering these questions requires knowledge on current affairs, corporate activities, and/or social issues. Although LLMs are successfully applied to various language tasks, models are pretrained on datasets consisting primarily of general English data. Consequently, their performance in other languages, such as Japanese, remains underexplored, particularly in domain-specific tasks.
Our study aligns with two existing directions: language-specific and domain-specific LLMs. Language-specific pretrained models have been developed for various languages (Zhang et al., 2021; dbmdz, 2023; tokyotech llm, 2023) nowadays, whereas domain-specific models excel in several domains such as finance (Scao et al., 2023). The combination of these two aspects with language models have been known promising since the era of BERTology (Beltagy et al., 2019; Ishigaki et al., 2023), however, many non-English languages, including Japanese, have no such LLM combinations for the recent decoder-only architecture such as GPTs (Brown et al., 2020). In this situation, first and foremost, a fundamental decoder-only model is required to facilitate future research to discover better pretraining methods. Therefore, as a case study, we developed the first Japanese LLM with 13 billion parameters for the business domain.
To leverage our Japanese business specific LLM in practice, it should be updated, for example, to answer the latest business-related questions. Thus, we adopt a pipeline approach for training our LLM: pretraining and continual updating. For the pretraining step, we train the model from scratch. Although there are two potential approaches – pretraining from scratch and continual pretraining from an existing LLM – we opt for the former because ample data are available in the target domain and language.
For the training dataset, we compile a diverse set that combines domain-specific and general-domain data. Specifically, we use over 40 billion tokens from patents and business documents to capture technical terminology, along with over 150 billion tokens from general-domain sources, such as Wikipedia, CC100, mC4, and Common Crawl.
To update our model with the latest business knowledge, we continually pretrain the model on the latest business documents collected in the past two months, ensuring that its knowledge remains current.
In addition to the commonly used general-domain benchmark such as ‘lm-evaluation-harness‘ (Gao et al., 2023), a business-specific benchmark is required to evaluate our business-specific model. For this purpose, we design a new dataset of 50 business questions for three different QA tasks: question-only, question with automatically retrieved context, and question with manually retrieved context. Manual human evaluations of the domain-specific benchmark demonstrate that our pretrained models outperform existing Japanese LLMs in terms of accuracy, particularly in the question-only setting. Moreover, we find that the updated models with the latest knowledge are more accurate than the pretrained models in answering questions about the most recent events in the last two months.
Our contributions are threefold: 1) we pretrain the first and largest Japanese domain-specific LLM on a new dataset; 2) we create a new benchmark with 50 questions for three different tasks; and 3) we demonstrate the effectiveness of our pretrained model in domain-specific tasks. This study establishes a fundamental environment for comparing future language- and domain-speciifc LLMs. Analysis presented in this paper provides insights for researchers who train domain- and language-specific LLMs in other domains and languages. Our model and benchmark questions are publicly available.  [^2] https://huggingface.co/stockmark.
Category Question
Current Affairs
Which country joined NATO in April 2023 in response to
Russia’s invasion of Ukraine?
Corporate Activities
Which Japanese startup has been developing perovskite
solar cells since 2022?
Social Issues
What is carbon neutrality?
Trends
What is a dark store?
Table 1: Examples of the business related questions translated from the original Japanese.
2 Methods
Our approach consists of three steps: collecting datasets, filtering them, and pretraining a LLM from scratch.
2.1 Dataset
To construct our pretraining dataset, we gathered 19.8% domain-specific and 80.2% general domain texts in Japanese. For the domain-specific data, we use business-related web pages and patent documents. The former, called the “Curated Business Corpus,” comprises publicly available web pages published by September 2023. We identified pages related to business using predefined URLs and a cue word list, and extracted the pages of which match the pattern of the URL, or which contain at least one of these keywords. The URLs and cue words are chosen to cover various aspects of the business domain, including chemistry, materials, biology, engineering, economics, current affairs, and social trends. In addition, we incorporate patent documents from the Japan Patent Office to enhance the model’s understanding of technical terminology. We collected as many articles as possible from these sources.
To ensure diversity of the training data, we include broader general-domain data from sources such as Wikipedia, CC100, mC4, and Common Crawl. These datasets contribute to general knowledge, which is essential for diverse natural language processing tasks.
2.2 Filtering
Filtering is critical for enhancing the dataset quality. We applied a three-step pipeline: language identification, removal of noisy characters, and deduplication.
Language Identification: Language identification is an important step in language-specific settings. To identify non-Japanese documents, we used a pipeline of two methods: the library-based and language characteristics-based. We initially used the “xlm-roberta-base-language-detection” library  [^3] https://huggingface.co/papluca/xlm-roberta-base-language-detection. Consequently, for documents with uncertain language identification results from the xlm-roberta-based tool, we applied the “franc” library  [^4] https://github.com/wooorm/franc. Any texts not recognized as Japanese were subsequently removed from the dataset.
Noise Character Removal: We believe that low-quality Japanese texts often take the non-sentential forms. Noisy texts are often composed only of dates, menu bars, URLs, or do not contain an end-of-sentence punctuation mark, “。” (period in Japanese.). Thus, we removed those texts from the corpus if they are regarded as non-sentential. This ensures the quality of our dataset and focuses on the model’s attention to relevant linguistic features. Also, because English sentences end with a period (“.”), using puctuation marks as a clue to identify non-sentential texts can aid in refining English sentences. Our characteristics-based method is simple but it has wide applicability in distinguishing languages that have specific characteristics from English; for example, Thai and Chinese do not have punctuation.
Deduplication: Finally, we deduplicated the collected documents and sentences to remove identical entries. For the document-level, we used “hash“, which is a built-in function of python, to convert documents into hashed values and removed duplicates memory-efficiently. For the sentence-level, we count the times of the sentence appearance and removed the sentences that appear more than 15 times. In both, document- an sentence-wise deduplication, we used exact matching of entries.
Following these preprocesses, our dataset comprised a total of 220 billion tokens, as shown in Table 2.
2.3 Pretraining
We use the same architecture as Llama2 (Touvron et al., 2023) with 13 billion parameters. We aligned our hyperparameters with those reported in the paper of Llama2.
For the training data in an epoch, we adopted the weighting strategy of the data used by Llama2, and doubled the amount of Wikipedia for two main reasons. First, Wikipedia data tends to be relatively clean. Second, Wikipedia contains a wealth of contents related to the business domain. In addition, we also double the data size of our Curated Business Corpus to enhance the infusion of domain-specific knowledge. By increasing the proportion of clean domain-specific data, we aim to mitigate the impact of noisy data sources such as mC4 and Common Crawl.
For the infrastructure, we used AWS’s Trainium, a hardware accelerator designed specifically for high-performance machine learning computations. We deployed our scripts on 16 trn1.32xlarge instances, each equipped with Trainium, to facilitate a robust distributed learning environment. The distributed learning process on Trainium was managed using the neuronx-nemo-megatron library  [^5] https://github.com/aws-neuron/neuronx-nemo-megatron, which is available on AWS’s custom accelerators. The pretraining phase spanned 30 days to complete training one epoch of the data.
Dataset Number of tokens [billion]
Curated business corpus 9.1
Patent 34.8
Wikipedia 1.0
CC100 10.9
mC4 53.2
Common Crawl 112.9
Table 2: The size of preprocessed dataset used for pretraining.
Dataset Number of examples language
dolly 15,015 translated Japanese
oasst 88,838 translated Japanese
alpaca 51,716 translated Japanese
Ichikara 10,329 Human-authored Japanese
Table 3: Candidates of datasets for instruction tuning.
2.4 Updating Pretrained Model with Latest Business Documents
In real-world applications, such as domain-specific QA, LLMs must be updated using the latest knowledge. However, the pretrained model discussed in the previous sections is trained on texts published only up to September 2023. It may be difficult to answer questions regarding events after October 2023.
To address this, we continue to train our pretrained model using the latest business documents published in October and November 2023 in our experiments. However, we are aware of the phenomenon of “catastrophic forgetting,” which occurs when acquiring new knowledge and potentially displaces previously learned information during pretraining.
To prevent this, we use a strategy that blends the latest business documents with randomly selected non-latest documents from the Curated Business Corpus. This strategy is inspired by previous research (Scialom et al., 2022). We introduce a hyperparameter $r$, which represents the proportion of the instances sampled from the set of non-latest documents; that is, we used the Curated Business Corpus as the source. For example, if $r$ is set to $0.3$, it means that 30% of the data of the continual update comes from the Curated Business Corpus. We investigate how different values of $r$ influence the performance of the domain-specific QA, and analyze the occurrence of catastrophic forgetting.
3 Experiments
To assess our LLMs, we create a benchmark for QA within the business domain. We then compared our models against different Japanese LLMs using this benchmark, along with the common Japanese benchmarks.
3.1 Benchmarks
Business Question Benchmark: This dataset comprises of 50 questions written in natural sentences, each paired with related articles. The questions cover various topics, including recent events, company activities, social issues, and business trends. Each question is linked to web pages retrieved through different methods: automatic and manual. The benchmark offers three QA settings: 1) answering without any context (NoContext-QA), 2) using an automatically retrieved webpage (AutoRAG-QA), and 3) using a manually retrieved webpage (ManualRAG-QA).
For the first task (NoContext-QA), web pages are not provided as a context. This setting allows us to assess the LLMs’ ability to generate answers based on the knowledge they contain internally, without knowledge existing outside the LLM.
For the latter two tasks (AutoRAG-QA and ManualRAG-QA), we extract only the text from each web page and truncate it to 1000 characters.
For AutoRAG-QA, we use a search engine to identify the most relevant web pages for each question. Subsequently, we selected the highest-ranked page of which the body text was available to be extracted as the context. However, because these web pages are automatically retrieved, they may not always provide answers to questions. Thus, this setting facilitates the assessment of how well LLMs can generate an answer, considering both the knowledge they already possess and any potentially irrelevant information on the web page.
For ManualRAG-QA we manually select a web page containing the answers to the question. Note that these web pages must contain answers to the questions; thus, this task is rather a comprehension task that requires better language reading ability than the ability to produce an answer from the knowledge in LLMs. We used the prompts shown in Appendix 1.
During the evaluation using this benchmark, the responses generated by the LLMs were manually assessed by an NLP researcher. The evaluation involves a binary judgement of the responses as either correct or incorrect, based on two main criteria. 1) Content faithfulness: the response must contain an answer to the question without any factual errors. 2) Response appropriateness: if the question included specific instructions (e.g. “provide only one example”), the response is considered correct only if it follows the instructions. The evaluator judges the output as correct if it satisfies both criteria. We used manual evaluations over automatic evaluation metrics because the latter (e.g., BLEU (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), BERTScore (Zhang et al., 2020)) do not adequately assess factual correctness, which is crucial in the business domain. Moreover, if a response contains repetition, for example, repeating the same sentence twice, redundant parts were disregarded, and the evaluation focuses solely on the unique content.
lm-evaluation-harness: lm-evaluation-harness  [^6] https://github.com/Stability-AI/lm-evaluation-harness is a common existing framework for assessing Japanese LLMs. It comprises eight Japanese NLP tasks in the general-domain. Each task is designed to probe different facets of language comprehension and generation. The tasks in this framework include reading comprehension (JSQuAD (Kurihara et al., 2022)), QA (JCommonSenseQA (Kurihara et al., 2022) and JAQKET  [^7] https://www.nlp.ecei.tohoku.ac.jp/projects/jaqket/), natural language inference (JNLI (Kurihara et al., 2022)), summarization (XLSum-ja  [^8] https://huggingface.co/datasets/mkshing/xlsum_ja), co-reference resolution (WinoGrande-ja (Takahashi et al., 2023)), and math problems (MGSM (Zhang et al., 2023)).
Owing to the unavailability of the MARC-ja dataset (Kurihara et al., 2022), we conducted lm-evaluation-harness on the remaining seven tasks. It is well acknowledged that the efficacy of LLMs is significantly influenced by the prompts or choice of the templates used in the evaluation. To this end, because multiple templates are available for a task, we conducted evaluations across all available templates, and the highest score is reported in this paper. The configurations of templates used in this study are listed in Table 9 for reproducibility.
3.2 Compared Models and Instruction-tuning
We compare our pretrained model with six Japanese LLMs and three multilingual LLMs. Japanese LLMs are categorized into two: models pretrained from scratch and models with continual pretraining from a multilingual model. The former, that is from scratch, is the same setting as ours, and three out of the six models are used for the comparison; 1) llm-jp-13b [^9] https://huggingface.co/llm-jp/llm-jp-13b-v1.0 is pretrained on the Japanese Wikipedia, mC4 and English Wikipedia, The Pile (Gao et al., 2020), and The Stack (Kocetkov et al., 2022), 2) plamo-13b [^10] https://huggingface.co/pfnet/plamo-13b is pretrained on English texts from RedPajama and Japanese texts from mC4 and the Japanese Wikipedia, and 3) weblab-10b [^11] https://huggingface.co/matsuo-lab/weblab-10b is pretrained on Japanese texts from mC4 and English texts from The Pile. The first two are models with 13-billion parameters and the last has 10 billion parameters.
We also compare our models trained by continual pretraining from existing multilingual models; 4) nekomata-14b [^12] https://huggingface.co/rinna/nekomata-14b is trained from qwen-14b on the Japanese CC100, C4, Wikipedia, Oscar, and Rinna curated corpus, and The Pile, 5) ELYZA-japanese-13b [^13] https://huggingface.co/elyza/ELYZA-japanese-Llama-2-13b is trained from Llama-2-13b on Japanese texts from Oscar and Wikipedia, 6) Swallow-13b [^14] https://huggingface.co/tokyotech-llm/Swallow-13b-hf is trained from Llama-2-13b on Japanese texts from Wikipedia and meticulously cleaned Common Crawl and English texts from RefinedWeb (Penedo et al., 2023) and The Pile. The model 4) has 14 billion while others have 13 billion parameters.
We also use multilingual LLMs for comparison; for 7) gpt-3.5-turbo-0125, 8) gpt-4-1106-preview, models are selected from the OpenAI API. 9) Llama-2-13b-hf (Touvron et al., 2023) [^15] https://huggingface.co/elyza/ELYZA-japanese-Llama-2-13b is a 13-billion English LLM, which can produce Japanese texts.
LLMs without instruction tuning rarely provide accurate answers to these questions, which suggests the necessity of this tuning. There are multiple data choices for instruction tuning in Japanese, such as Ichikara, Alpaca, Dolly, and Oast. The statistics for these datasets are presented in Table 3. For the dataset selection, we conduct preliminary experiments using our proposed model detailed in Table 4. The values in the table demonstrate that all instruction-tuned models generally perform similarly, with average scores of approximately 0.550, except for the model trained on the Oast dataset, which achieves only 0.510. However, the model tuned on the Ichikara dataset consistently performs well on question-answering tasks, achieving scores of 0.78 on JSQuAD, 0.86 on JAQKET, and 0.84 on JCommonsenseQA. Therefore, we opt for the Ichikara dataset for instruction tuning.
Note that both our LLMs and the compared models are instruction-tuned on Ichikara dataset  [^16] https://liat-aip.sakura.ne.jp/wp/llmのための日本語インストラクションデータ作成/llmのための日本語インストラクションデータ-公開/. We use Low-Rank Adaptation (LoRA) Hu et al. (2021) for the instruction tuning.
model JSQUaD JAQKET JCom. XW. JNLI MGSM XLSum Ave.
Ichikara 0.78 0.85 0.84 0.75 0.49 0.08 0.08 0.550
Alpaca 0.73 0.85 0.81 0.73 0.57 0.07 0.07 0.545
Dolly 0.77 0.87 0.81 0.73 0.53 0.08 0.08 0.547
Oast 0.77 0.85 0.81 0.74 0.25 0.07 0.08 0.510
Table 4: Preliminary experiments: the comparison between datasets for instruction tuning. All base-model is our proposed 13-billion Japanese business domain specific LLM. JCom. and XW. refer to the JCommonSense and XWinograd datasets, respectively. The Ave. column shows the averaged values over all datasets.
4 Results
Results on Domain-specific Benchmark: Table 5 presents the results on the domain-specific benchmark. The table is divided into three categories representing different models: Japanese LLMs pretrained from scratch, Japanese LLMs with continual pretraining, and multilingual LLMs.
In the case of NoContext-QA, our pretrained model achieves the highest accuracy at $0.90$, compared to other existing Japanese LLMs such as nekomata-14b and Swallow-13b, which score $0.78$ for both. In this setting, the LLMs must generate answers from internally stored knowledge without using external knowledge such as retrieved documents. This suggests that our model possesses domain-specific knowledge within LLM.
For ManualRAG-QA, our model achieves the best score among the LLMs pretrained from scratch. For example, our model achieves an accuracy of $0.84$, whereas the other full-scratch models achieve accuracies of $0.76$, $0.82$, and $0.84$, respectively. Similar tendencies are observed in the AutoRAG-QA task. Our model achieves $0.74$ whereas the other models achieve only values of $0.62$, $0.64$, and $0.34$, respectively. These RAG tasks involve comprehension and rely on language ability to find answers in documents. This is in constrast to NoContext-QA, which tests the model’s ability to generate answers without document retrieval. From these results, we conclude that our pretrained model excels in answering business-specific questions because of its knowledge; however, the Japanese LLMs with continual training can achieve better performance in language comprehension tasks.
Comparing the two strategies for pretraining (from scratch or continual), continual pertinence consistently outperforms the models pretrained from scratch. Among the general-domain Japanese LLMs, the best performing LLMs with continual pretraining i.e., Swallow-13b and nekomata-14b, achieve $0.78$ in NoContext-QA, whereas the best general-domain LLM pretrained from scratch i.e., llm-jp-13b, achieves a notably lower score of $0.34$. Despite the fact that the full scratch strategy seems disadvantageous overall, it should be noted that our model performs the best on the NoContext-QA task among all compared models, which supports the effectiveness of the use of domain-specific data in pretraining. For the QAs with contexts i.e., ManualRAG-QA and AutoRAG-QA, our model achieves the best score ($0.84$ and $0.74$) among the full scratch models but the best performing models with continual pretraining perform better ($0.94$ and $0.82$). One of possible reasons is that continual models e.g., Swallow-13b or ELYZA-japanese-13b, use Llama2 as a base model, which is pretrained on a larger dataset. Thus, the base model might provide a strong language comprehension ability. Although this paper focuses on pretraining models from scratch, continually learning a strong existing model such as Llama2 on our domain-specific corpus may produce much more performance. We leave further investigations for future work.
model NoContext-QA ManualRAG-QA AutoRAG-QA
Japanese LLMs trained from scratch
-   Our model 0.90 0.84 0.74
- 1) llm-jp-13b 0.34 0.76 0.62
- 2) plamo-13b 0.34 0.82 0.64
- 3) weblab-10b 0.26 0.52 0.34
Japanese LLMs with continual pretraining
- 4) nekomata-14b 0.78 0.74 0.76
- 5) ELYZA-japanese-13b 0.32 0.94 0.70
- 6) Swallow-13b 0.78 0.90 0.82
Multilingual LLMs
- 7) gpt-3.5-turbo-0125 0.54 0.62 0.34
- 8) gpt-4-1106-preview 0.78 0.94 0.86
- 9) Llama-2-13b-hf 0.24 0.84 0.64
Table 5: Results on business-specific benchmark. Each score stands for the accuracy. The values obtained from the best performing models in each category is shown in bold.
Results on General-domain Benchmark: Table 6 presents the results on the general-domain benchmark, “lm-evaluation-harness.” Among the models pretrained from scratch, our model demonstrates superior performance in terms of the averaged score i.e., ours achieves $0.55$ on average while the best performing Japanese LLMs pretrained from scratch achieve $0.49$. In particular, in QA tasks without context, such as JAQKET and JCommonSense, our model achieves the best scores $0.78$ and $0.85$, respectively.
model jSQUaD JAQKET
JCom.
XW. JNLI MGSM XLSum Ave.
Japanese LLMs Pretrained from Scratch
- Ours 0.78 0.85 0.84 0.75 0.49 0.08 0.08 0.55
- 1) weblab-10b 0.72 0.43 0.65 0.67 0.30 0.02 0.05 0.41
- 2) plamo-13b 0.68 0.69 0.64 0.68 0.41 0.02 0.10 0.46
- 3) llm-jp-13b 0.69 0.76 0.79 0.70 0.37 0.02 0.10 0.49
Japanese LLMs with Continual Pretraining
- 4) nekomata-14b 0.87 0.88 0.94 0.80 0.65 0.36 0.22 0.68
- 5) ELYZA-13b 0.79 0.75 0.87 0.78 0.51 0.10 0.19 0.57
- 6) Swallow-13b 0.86 0.91 0.91 0.72 0.52 0.18 0.20 0.61
Multilingual LLMs pretrained from Scratch
- 9) Llama-2-13b-hf 0.81 0.75 0.82 0.63 0.47 0.12 0.21 0.54
Table 6: Results on lm-evaluation-harness. JCom. and XW. refer to the JCommonSense and XWinograd datasets, respectively. The Ave. column shows the averaged values over all datasets. OpenAI’s GPTs cannot be used for this experiments because the model parameters are not public.
Analysis of Models with Latest Business Documents: We compare the updated model with the original pretrained one. The left part of Table 7 presents the validation loss values on two datasets: the 8,192 documents used for the validation portion in pretraining (Pretrain Data) and the latest business documents used for updating the model (Latest Data). The loss value of the updated model with $r=0.0$, indicating it is trained only on the latest documents, is 2.05 on the latest data. This is an improvement over the original pretrained model’s $2.25$. This indicates that the updated model better fits the latest documents. However, the loss on the pretaining data worsens from 2.11 to 2.19, compared to the original model. Increasing the value of $r$, that is adding more pretraining data to the latest documents, can alleviate the increase in the loss of the pretraining data. When we increased $r$ to $0.3$, we achieve a loss of 2.12, nearly equivalent to the original model’s 2.11, without degrading the loss on the latest documents (2.04).
To investigate the capability of our model to answer questions that require the latest knowledge. We create a set of 10 questions about latest business topics, specifically from October to November 2023. We selected target topics that showed a notable increases in access to a search engine, compared to that in September 2023. The accuracy of NoContext-QA on the non-latest questions and latest questions is shown in the right part of Table 7. In the latest questions, the updated models demonstrate higher score ($0.90$) than the pretrained model ($0.30$), denoting that a Japanese- and business-specific model can learn new information with the continual learning. Different from the results in terms of loss, we did not observe significant changes in terms of accuracy among the models with different values of $r$.
Loss $\downarrow$ Accuracy $\uparrow$
Pretrain Data Latest Data Non-LatestQ. LatestQ.
Pretrained model 2.11 2.25 0.90 0.30
Updated model ($r=0.0$) 2.19 2.05 0.92 0.90
Updated model ($r=0.1$) 2.14 2.05 0.92 0.90
Updated model ($r=0.3$) 2.12 2.04 0.90 0.90
Table 7: Comparisons of validation losses on the pretrain data and the latest data, and accuracies on the questions about the latest business and non-latest business knowledge (Latest Q. and Non-latest Q.).
Examples of Outputs Table 8 shows the examples of the outputs from our pretrained model and compared LLMs. When we input a question “Which two banks failed in March 2023?”, our pretrained model correctly generates two banks: the Silicon Valley Bank and the Signature Bank. Other three Japanese LLMs’ outputs contain hallucinations, which are indicated in bold in the table. For example, Wells Forgo, Bank of America, and Citigroup did not fail in March 2023. We observed a similar tendencies for the other questions.
Model Generated Text
Ours The two U.S. banks that failed in March 2023 were Silicon Valley Bank and Signature Bank. (2023年3月に経営破綻した米国の2 つの銀行は、シリコンバレー銀行とシグネチャー銀行である。)
llm-jp-13b-v1.0 The two U.S. banks that failed in March 2023 were Wells Fargo and Bank of America. (2023年3月に経営破綻したアメリカの2つの銀行は、ウェルズ・ファーゴとバンク・オブ・アメリカである。
plamo-13b The two banks that failed in March 2023 were Bank of America and Citigroup. (2023年3月に経営破綻した2つの銀行は、バンク・オブ・アメリカとシティ・グループでした。)
weblab-10b The two U.S. banks that failed in March 2023 were Bank of America and Wells Fargo. (2023年3月に経営破綻したアメリカの2つの銀行は、Bank of America と Wells Fargoです。)
ChatGPT (gpt-3.5-turbo-0125) The two American banks that went bankrupt in March 2023 are Regional Asahi Bank and Centry National Bank. (2023年3月に経営破綻したアメリカの2つの銀行は、リージョナル朝陽銀行とセントリーナショナル銀行です。)
Table 8: Examples of output generated from different LLMs. The sentences in the buckets are the original outputs in Japanese. Hallucinated terms are presented in bold.
5 Related Work
Recent major LLMs are multilingual models such as Llama2 (Touvron et al., 2023), OpenAI’s GPTs (OpenAI, ), and BLOOM (Scao et al., 2023). The datasets used for such major models often include a high percentage of English data. Therefore, performances in languages other than English have been still underexplored. Studies targeting non-English languages, such as Chinese (Zhang et al., 2021), German (dbmdz, 2023), and Japanese (tokyotech llm, 2023), increase the proportion of non-English texts used in pretraining. Whereas, we propose an approach to build a large size of Japanese specific corpus by filtering out other langauges with language identification libraries and a noise character detection, and we pretrain Japanese-specific LLM.
The language-specific datasets for instruction tuning have been released. For Japanese, Ichikara, Alpaca, Dolly, and Oast are common; however, the performances of models trained on these datasets have not been compared in depth. Our preliminary experiment is the first to compare these datasets, which can be considered as an important contribution for the Japanese LLM community. Also, our experiments can provide insights for researchers who focus on other non-English languages.
Domain-specific models are pretrained on two different ways; training from scratch or continual training. Representative examples of the former started for the encoder-only models such as SciBERT (Beltagy et al., 2019) in the science domain and later decoder-only large language models follow, e.g., BloombergGPT (Wu et al., 2023). The continually learned domain-specific LMs include BioBERT (Lee et al., 2019), exBERT (Wang et al., 2020). Scratch approaches are often been used for settings where we can get sufficient data. Our setting is categorized into this setting, thus, we use the scratch approach.
The combination of the two directions, i.e., language- and domain-specific settings, is underexplored in Japanese. Japanese domain-specific language model does not exist, except for the pretrained BERT for the material science domain (Ishigaki et al., 2023). Our target is decoder-only architecture with many more parameters for the business domain, which has high demand in the industry but is less studied. Our pretrained model is the first Llama-based domain-specific model for Japanese
Continual pretraining is a promising direction if we have only a small dataset for pretraining. Our experiments for updating the pretrained model are categorized in this setting. “Catastrophic forgetting” is a major problem in this case (Ling et al., 2023). Scialom et al. (2022) suggest that the mixture of two types of data can mitigate this problem, thus, we used the technique to mix the latest documents and older ones.
6 Conclusion
This paper presented the first Japanese business domain-specific LLM. We pretrain the LLM with 13 billion parameters from scratch and will be made public. Evaluations demonstrated that our LLM achieves better on the domain-specific benchmark we newly released. We also found that updating the pretrained parameters with the latest business documents can make the model better fit the latest documents. For future work, we will explore methods that leverage the context in RAG setting while utilizing the knowledge within LLMs.
7 Ethics Statement
The proposed model is still in the early stages of research and development, and its output has not yet been adjusted to align with safety considerations. However, adjustments will be made in the future to ensure that the model takes safety into account.
Appendix A Appendix
A.1 Templates used in Experiments
Table 9 shows the versions of templates used in our experiments.
Task Task Version Prompt Version Number of Few-shot
JSQuAD 1.1 0.1, 0.2, 0.3 2
JAQKET v2-0.2 0.1, 0.2, 0.3 1
JCommonsenseQA 1.1 0.1, 0.2.1, 0.3 3
JWinograd ja - 0
JNLI 1.3 0.2, 0.3 3
MGSM 1.0 0.0, 0.3 5
XLSum ja-1.0 0.0, 0.3 1
Table 9: Examples of output generated from different LLMs. We show the English translations while the sentences in the buckets are the original outputs in Japanese. Hallucinated terms are in bold.
A.2 Prompts used in Experiments
Please answer the question briefly. Question:{question} ### Output:
Figure 1: Translated prompt for NoContext-QA. In the experiment, Japanese prompt was used.
Please answer to the given question. If the answer to the question is included in the article text, please use the answer from the text. If the article does not contain the answer please state that ”the article does not contain the answer” and answer the question using your knowledge. Question:{question} Article Text: {the first 1k characters} ### Output:
Figure 2: Translated prompt for ManualRAG-QA and AutoRAG-QA. In the experiment, Japanese prompt was used.